NPP Projection Report 8/13 Open/Mid


To do this evening/ was completed: Weekly update is sent out, please give it a read.

Presentation Issues:

Preshow Clips/Playlist:

Trailer Updates:

Special Events/ AV set up: Ran loops for Bad Boys.

I programmed the remaining spot settings for Showgirls. I wasn't sure if it was the orange and spots or red/orange and spots that was needed so I programmed both. Didn't look like either were saved anywhere.

Realm of Satan went well. I recorded the Q&A but the internet on the T2 PC is acting up. Will upload it tomorrow.

Ordering/Receiving (posters & trailers): 

Ordering/Receiving (content): Ingested Caught By The Tides

Outgoing/Shipping (content):

Print/video Inspection:

Asset tracking doc updated/checked?:

Maintenance/cleaning:

Cue sheets updated/checked (every shift):

Marquee updated?:

Posters updated?:

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Open 12/28

NPP Projection Admin (Lillian/Will) 7/7

Projection Admin 9/8 Lillian/Will